Break-Even Sales and Cost-Volume-Profit Chart

Last year, Hever Inc. had sales of $500,000, based on a unit selling price of $250. The variable costper unit was $175, and fixed costs were $75,000. The maximum sales within Hever Inc.'s relevant range are 2,500 units. Hever Inc. is considering a proposal to spend an additional $33,750 on billboard advertising during the current year in an attempt to increase sales and utilize unused capacity.

Required:

1. Construct a cost-volume-profit chart on your own paper, indicating the break-even sales for last year.

Break-even sales (dollars) _________________
Break-even sales (units) ___________________

2. Using the cost-volume-profit chart prepared in part (1), determine (a) the income from operations for last year and (b) the maximum income from operations that could have been realized during the year.

Income from operations ________________
Maximum income from operations __________________

3. Construct a cost-volume-profit chart (on your own paper) indicating the break-even sales for the current year, assuming that a noncancelable contract is signed for the additional billboard advertising. No changes are expected in the unit selling price or other costs.

Dollars ______________
Units _______________

4. Using the cost-volume-profit chart prepared in part (3), determine (a) the income from operations if sales total 2,000 units and (b) the maximum income from operations that could be realized during the year.

Income from operations at 2,000 units ______________
Maximum income from operations ______________________
Answer #1

1. Construct a cost-volume-profit chart on your own paper, indicating the break-even sales for last year.

Contribution margin = unit selling price - variable costper unit

Contribution margin =(250-175)

Contribution margin = 75

Contribution margin Ratio = Contribution margin /unit selling price

Contribution margin Ratio = 75/250

Contribution margin Ratio = 30%

Break-even sales (dollars) =  fixed costs /Contribution margin Ratio

Break-even sales (dollars) = 75000/30%

Break-even sales (dollars) = 250000

Break-even sales (units) = fixed costs /Contribution margin

Break-even sales (units) = 75000/75

Break-even sales (units) = 1000

Answer

Break-even sales (dollars) $ 250,000
Break-even sales (units) 1000

2. Using the cost-volume-profit chart prepared in part (1), determine (a) the income from operations for last year and (b) the maximum income from operations that could have been realized during the year.

No of Unit sold = sale /Sale Price = 500000/250 = 2000

Income from operations for last year = Contribution margin*No of Unit sold - Fixed cost

Income from operations for last year = 75*2000 - 75000

Income from operations for last year = $ 75000

Maximum income from operations = Contribution margin*No of Maximum Unit can be sold - Fixed cost

Maximum income from operations = 75*2500 - 75000

Maximum income from operations = $ 112500

Income from operations $ 75000
Maximum income from operations $ 112500

3. Construct a cost-volume-profit chart (on your own paper) indicating the break-even sales for the current year, assuming that a noncancelable contract is signed for the additional billboard advertising. No changes are expected in the unit selling price or other costs.

Contribution margin = unit selling price - variable costper unit

Contribution margin =(250-175)

Contribution margin = 75

Contribution margin Ratio = Contribution margin /unit selling price

Contribution margin Ratio = 75/250

Contribution margin Ratio = 30%

Total fixed costs = 75000+33750 = 108750

Break-even sales (dollars) =  fixed costs /Contribution margin Ratio

Break-even sales (dollars) = 108750/30%

Break-even sales (dollars) =362500

Break-even sales (units) = fixed costs /Contribution margin

Break-even sales (units) = 108750/75

Break-even sales (units) = 1450

Dollars $ 362,500
Units 1450

4. Using the cost-volume-profit chart prepared in part (3), determine (a) the income from operations if sales total 2,000 units and (b) the maximum income from operations that could be realized during the year.

No of Unit sold = sale /Sale Price = 500000/250 = 2000

Income from operations for last year = Contribution margin*No of Unit sold - Fixed cost

Income from operations for last year = 75*2000 - 108750

Income from operations for last year = $ 41250

Maximum income from operations = Contribution margin*No of Maximum Unit can be sold - Fixed cost

Maximum income from operations = 75*2500 -108750

Maximum income from operations = $ 78750

Income from operations at 2,000 units $ 41250
Maximum income from operations $ 78750